leaqueen:

The table below shows the size of nine families selected at random from two neighborhoods in a large city: Family Size (in number of people) Neighborhood Q 2 5 4 3 2 5 3 6 5 Neighborhood S 2 3 2 3 7 2 3 3 2 Which neighborhood appears to have a bigger family size? Explain your answer using two or three sentences. Make sure to use facts to support your answer.

IhelpUuHELPme:

Well, to find the neighborhood that has the bigger family size, you would have to add the number of families in each neighborhood. Q = (2 + 5 + 4 + 3 + 2 + 5 + 3 + 6 +5 = ?) S = (2 + 3 + 2 + 3 + 7 + 2 + 3 + 3 + 2 = ?) Once you add those numbers for each neighborhood, the one with the higher number is the one with the bigger family size. Correct me if I'm wrong xd

leaqueen:

omg thx